OpenStudy (anonymous):

The length of a rectangle is 8 in. more than its width. The perimeter is 52 in. What is the length of the rectangle?

Nnesha (nnesha):

do you know the perimeter equation

Nnesha (nnesha):

it is p= 2w + 2l w is width and l is length

Nnesha (nnesha):

The length of a rectangle is 8 in. more than its width. now translate this into the equation Length = l is mean equal sign more than = plus width = w so l = 8 + w

Nnesha (nnesha):

and perimeter is 52 so we can replace p by its value 2w + 2l = 52 now l= 8 + w is already solved for l so plug l value in to 2w + 2l=52 this equation where l is

Nnesha (nnesha):

 now distribute bracket by 2